### v2

# Correlation studies of the 7He excited states

Abstract
The unbound nucleus 7He was recently studied in the 2H(6He,1H)7He reaction at 29 MeV beam energy in Ref. [M. S. Golovkov et al., Phys. Rev. C 109, L061602 (2024)]. The excitation spectrum of 7He was measured up to MeV ( is energy above the 6He- threshold). Angular distribution for the 6He- decay of the 7He ground state can be explained by a strong spin alignment induced by a reaction mechanism. The correlation information for the higher-lying 7He excitations is available as backward-forward asymmetry for the 6He- decay in the 7He frame. The asymmetry function has an expressed energy profile which may be explained by using quite restrictive assumptions about structure of 7He excitations or/and peculiarities of the reaction mechanism. In the analysis of [M. S. Golovkov et al., Phys. Rev. C 109, L061602 (2024)] the observation the state in 7He is declared with MeV. Our work is based on the same 7He data. However, the data analysis was improved and also the data interpretation is substantiated with the detailed PWBA reaction studies and coupled-channel calculations of the 7He continuous spectrum. The idea of the resonant state with is rejected. In addition, the position of the state in 7He is confined to the interval MeV, with “preferred” value 2.6 MeV. There is indication on the second state in the data with MeV and with the lower resonance energy limit MeV. Importance and prospects of more detailed correlation studies of 7He continuum are discussed.
I Introduction
Clear understanding of excitation spectra of the lightest nuclei is critical for general understanding of nuclear structure and nucleon-nucleon interaction in nuclei. The 7He is intriguing system in this sense. We know the ground state (g.s.) properties very well. The state at MeV above the 6He- threshold is clearly identified because of its unique dominating decay pattern 4He+ [1]. However, information about other 7He excitations is nebulous.
[…]
Generally, the 7He system has been studied many times and detailed reviews of this work can be found in [6] and [7]. There is the g.s. at MeV and some broad overlapping structures are typically observed above it. The evident candidate to be present in this structures is the spin-orbit partner of the ground state. The review [7] split the results in this field into two “camps”: (i) the works which support existence of the low-lying excited state with MeV in 7He [8, 9, 10, 11, 12, 13, 14, 15], and (ii) the works which are in favor of the with MeV [16, 17, 18, 19, 20, 3, 21, 22, 23, 24]. Such a dichotomy is a bit artificial, however it well demonstrates a strong disagreement among different studies, both theoretical and experimental. The 6He- correlation data obtained in this work indicate that the state should be reasonably low-lying ( MeV) and should be well populated (with population cross section, comparable with the cross section for the g.s.).
The g.s. is well known to have non-single-particle structure with important (even dominant) 6He+ wave function (WF) component. Qualitatively this means that the excited states represented by the orthogonal (to the g.s.) mixtures of the 6He(g.s.)+ and 6He+ WF components may be situated reasonably low in excitation energy. There are various theoretical calculations predicting the state of 7He at MeV [25, 24, 26]. There is indication on the second state in our 7He data with MeV and with the lower resonance energy limit MeV.
We also discuss possibility of the resonant state at about MeV as predicted in [26] and observation of such a state in 7He is declared in [5] with , MeV. The possibility of the state was for a long time discussed in 9He [27, 28, 29, 30]. From qualitative point of view the principal possibility of such a state in 9He is supported by evolution of the “intruder” orbital along the isotone: there are ground states in 10Li, 11Be and low-lying excitations in 12B, 13C build on the configuration. No such support for the low-lying state of 7He can be found in the isotone: only in 9Be there is a weak evidence for the state above the 8Be+ threshold. Theoretical works except the paper [26] all predict either no low-lying state or repulsion in the 6He+ channel [31, 24, 25, 32, 33, 34]. In our work we demonstrate that the on-shell R-matrix parameterization used in the work [5] to infer the resonance properties in 7He is not applicable for the reaction studies (where the off-shell T-matrix should be applied).
The situation when we do not understand such a basic structure characteristic as spin-split between spin-orbit partners and the ground state multiplet structure in such a “simple” light nuclide as 7He is quite unsatisfactory and it is calling to a dedicated research. So, the aim of this work is to get a deeper insight in these questions by using the correlation information obtained in the 2H(6He,1H)7He reaction.
The system of units is used in this work.
II Experimental setup and method
The experiment was carried out at the fragment separator ACCULINNA-2 [35] at U-400M heavy ion cyclotron, at Flerov Laboratory of Nuclear Reactions (JINR, Dubna), using the method of double -6He and triple -6He- coincidences, developed in [36].
II.1 Experimental setup
The consistant description of the experimental setup is presented in [4]. Here we provide only a brief account of the main parts of the experimental setup which we consider useful for the futher discussion. Fig. 1 shows schematic layout of the experimental setup with general dimensions.
[…]
The protons from the reaction were registered by four telescopes (-telescopes), each consisting of two layers of double-sided silicon strip detectors. The -telescopes were covered the 150∘–170∘ range of laboratory angles.
[…]
II.2 “Combined” mass
Important opportunity connected with neutron data is the complete-kinematics reconstruction of events. The registration of neutrons from the 7He decay provides information about its decay properties, in particular, the decay energy. The neutron-coincidence events formally contain redundant information about 7He decay energy and there exists an approach that allows to use this redundant information for improvement of the experimental resolution.
[…]
II.3 Correlation measurements
[…]
For the events with neutron coincidences (with complete kinematics) the corresponding correlation functions can be directly extracted from data. This analysis is discussed in the Sec. VII.2.
Because the experimental setup accepts only the events where the alignment direction is approximately opposite to the beam direction, this effect should show itself in the distribution of the heavy reaction fragment over the longitudinal velocity. Analysis of this data is discussed in the Sec. VII.1.
III Helium-7 excitation spectrum
With the experimental setup Fig. 1 a standard way to get the 7He spectrum is construction of its missing mass by using the information from the detection of a proton recoil. Figure 3 (a) shows the 7He MM augmented with information from the forward ToF detector — registration of 6He and 4He heavy products from the 7He decay, see the gray rectangle in Fig. 2. The ToF data here is recalculated in terms of the heavy fragment longitudinal momentum in the 7He frame. It can be seen in this plot that the MM background conditions are quite good in the experiment, with the major background contribution coming from the random coincidences with the 6He beam located at about MeV. The kinematical cut-off in this plot allows to drastically reduce the background in the MM 7He spectrum, see the result in Fig. 3 (b). It should be understood that 4He heavy products originating from the 7He + + decay channel are confined inside the kinematical locus shown in Fig. 3 (a), and, thus the MM spectrum of Fig. 3 (b) includes this channel as well.
The energy resolution of the 7He MM spectrum in Fig. 3 (b) was obtained as keV FWHM in the energy range MeV and keV FWHM in the energy range MeV by the Monte-Carlo simulations. The results of the simulations are well confirmed by the observed width of the 7He ground state peak. The 7He MM spectrum has a sharp cut-off at about MeV connected with threshold for registration of the slow protons in the backward telescopes. The results of the setup efficiency simulations are shown in Fig. 3 (b) by the blue hatched regions. The simulations are shown with arbitrary scaling and normalazed on peak values at 0.4 MeV. The lower bounds of these regions correspond to assumption of isotropic 7He + decay. The upper bounds correspond to the most extreme correlated decay case when the products are strongly focused in the forward/backward direction in the frame aligned with transferred momentum, see this discussion further in Sec. VII.2. This strong focusing is described by pure term in Eq. (25). Using the efficiency corrections we can find that from of the total cross section with MeV are connected with population of excited states. In the first instance this may be seen as evidence for low spectroscopy of the 6He g.s. configuration in 7He g.s. structure.
A more advanced treatment of the 7He is available using the information from neutron coincidence data. It can be seen in Fig. 3 (c) that efficiency of the neutron wall is something around on the 7He g.s. and it rapidly drops down at MeV rendering the spectrum practically nonexistent at MeV. As it was discussed above in Section II.3 the combined mass spectrum reconstruction appears to be drastically more precise compared to missing mass providing the keV FWHM energy resolution for the 7He g.s., see the red histogram in Fig. 3 (c).
The precise measurement of the 7He g.s. width may furnish additional information about 7He structure. The widely used nuclear levels compilation [43] provides MeV for the 7He g.s. In the further works [44, 45, 22, 46, 6] the estimates of are provided in the range MeV. It should be noted that none of these works have made a crucial effort to determine this value precisely and disagreement among different experimental values is typically larger than their declared errors. The combined mass method gives a high enough resolution for the 7He g.s. width estimation. Such kind of estimates have been given in [4]. In this paper we present the results of more enhanced analysis of the 7He g.s. parameters. The 7He combined mass spectrum is shown in Fig. 4 on a larger scale together with MC simulations for different 7He g.s. widths values. The analysis of the analysis, see the inset in Fig. 4, gives estimates of the resonance energy and width MeV. The errors (confidence intervals) we define by a criterion . One can see that the new results is quite close to the results of [4] ( and MeV). However the new estimate have greater value and it confidence interval is significantly wider, therefore it’s consistent with the results of the majority of previous works. We consider that obtained precision for value is mainly determined not by measurement method, but by the available statistics. According to our estimates, in a prospective experiment in analogous technique (with reasonable statistics) the precision 10–15 keV may be aimed.
IV PWBA model for the 2H(6He,1H)7He reaction
Important qualitative features of the 2H(6He,1H)7He reaction can be understood based on the the plane-wave Born approximation (PWBA) model, which well reflects the “quasifree scattering” aspect of the process. Within the historical terminology for this class of reactions we consider the system of three particles, which, in the anti-lab system, are “target” (6He), and composite “beam” (2H), consisting of “spectator” (proton), “participant” (neutron), see Fig. 5. The experiment is performed in the inverse kinematics and we keep “beam” and “target” notation with quotation marks for this Section to avoid confusion here. Completely analytical treatment of the model is possible under two essential physical assumptions:
Plane wave motion is assumed in the initial and final state coordinates .
Only one interaction is taken into account in the T-matrix — the potential between “target” and “participant” fragment.
There is some confusion in terminology, but it looks that the most widespread reading is that quasifree scattering (QFS) is further simplification of PWBA, when the off-shell T-matrix is replaced with the experimental or phenomenological cross section (on-shell T-matrix), sometimes with simple extrapolation in the off-shell region.
Since 1950s [47, 48] this approach was elaborated on numerous occasions [49, and Refs. therein]. It was especially popular for reactions with nucleon transfer from deuteron, having the low binging energy and huge geometric extent, which make the above physical approximations easy to justify. Nevertheless, we found it necessary to provide a detailed discussion of the model in this work for several reasons: (i) it is difficult to find in the literature consistent discussion of the method in the aspect concerning correlations in the decay of unstable product; (ii) the approach has never been discussed in sufficient details in our previous experimental works [39, 40, 29, 41, 42], (iii) the rather detailed discussion of correlations in this work would be hard to follow otherwise.
Figure 5 shows the Jacobi coordinates for prior and post forms of the wave function. These sets have a simple relations
[…]
the T-matrix has explicitly factorized form (the 6He “target” particle is spinless)
where the initial state effects are realized via formfactor , while the “target”-participant interaction properties are reflected in the off-shell T-matrix of the “quasifree subsystem”.
It should be emphasized that there are two transferred momenta in the model: (momentum transfer to spectator in the projectile rest frame) and (momentum transfer to the quasifree channel). These momenta are defined by relation between initial and final state Jacobi coordinates:
The behavior of the transferred momenta within the kinematical range of interest for the experimental conditions is illustrated in Fig. 6.
The transfer formfactor (the simple -wave deuteron WF is assumed) and the elastic scattering -matrix are defined in a standard way
The triple-differential cross section is
and five-fold-differential cross section suitable for 7He correlation studies can be presented in the density-matrix formulation, which is specially convenient for the further phenomenological studies
The density matrix depends on the excitation energy
and the center of mass reaction angle , associated with the momentum , while the transition amplitudes depend on and
The partial components of the elastic -matrix and transfer formfactor for the simple -wave only deuteron WF with radial component are defined as
The matrix is pure angular density matrix
with symmetries
It is normalized to give unity for each state
The maximum spin alignment in the “quasifree” 7He channel is realized in the system, where . In this frame the density matrix gets the most sparse and simple form. For the state vector one gets
The scattering WF diagonalizing the elastic -matrix, is normalized by the asymptotic condition
For that reason the PWBA transition amplitude can be represented as
where the are real-valued functions. For reactions well suited for the quasifree approximation, we may expect, although can not be completely confident, that the real transition amplitudes and phases are reasonably close to the PWBA values. The population probabilities for different states can be defined as
where phenomenological corrections, if necessary, are included via spectroscopic factors .
The correlations pattern for the state set in the frame where directed along axis can be obtained as (case of a complete equatorial alignment)
where the variable and the relative phases are
There are two important features clarified within the PWBA model at that point:
The reaction for our kinematical conditions are not “comfortable” for interpretation in terms of “quasifree scattering”. The values of and transferred momentum entering the are shown in Fig. 7. The off-shell effect is large in this reaction and this effect is strongly varying across the 7He excitation energy range MeV accessible in the experiment. According to Fig. 7 one may expect that quasifree scattering approximation may become safe (small off-shell corrections) at MeV.
[…]
V Models for the 7He continuum
The easiest approximation for the 7He continuum states are provided by the single-channel potential model. For simplicity we used the square-well potential, which is sufficient for qualitative considerations and allows to make most of calculations analytically.
V.1 The ground state
The 7He ground state width is around 140 keV. With potential of reasonable radius fm one get width of keV. To play with the width values in potential model we can vary the potential width. This is not quite consistent with physics of the case, where reduction on the 7He g.s. is connected with not single particle nature of this state (strong mixing with + configuration). The continuum profile calculated with fm and giving keV are shown in Fig. 8 (a). In this model the high-energy “potential tail” of the resonance is scaled as . This behavior is important for understanding of the state properties (discussed below) and correlations (see Section VII.1.2).
Within the QFS approximation to PWBA [4, 5] instead of the off-shell T-matrix the phenomenological on-shell cross section is used, which is extrapolated off-shell in some reasonable way. The on-shell population probability Eq. (12) is related to the elastic cross section as
From the T-matrix definition (7) a reasonable off-shell extrapolation is
see, e.g., [50, 49]. It is evidently precise in the low limit, which is not a good approximation for the experimental conditions, see Fig. 7. Nevertheless, it can be found in Fig. 8 (b) that the off-shell extrapolation is quite precise in the case of -wave states, represented by potential scattering.
V.2 The state
It is typically expected that interaction in this partial wave is quite featureless repulsion. The continuum profiles calculated with repulsive potential and deep attractive potential (with “forbidden state”) are shown in Fig. 9. The phase shifts for these cases are fine-tuned to be very close to each other and to the phase shifts obtained in sophisticated continuum shell-model calculations [24, 34]. One may see in Fig. 9 that the off-shell T-matrices are (i) strongly (factor ) different depending on specific dynamics and (ii) they are evidently not “off-shell extrapolatable” by simple expressions like Eq. (16). This is demonstration that treatment of the -wave populations in reactions should be considered with more caution. This fact is also important for discussion of possible interpretation of correlations in terms of -wave resonance contributions, see Sec. VII.1.3.
For attractive potential the PWBA -wave population is on the upper limit of what is admissible from experimental data. So, phenomenologically this population can not be increased, but can be reduced (e.g. repulsive potential). From correlations point of view if we reduce the -wave population of Fig. 9 50-fold, it will be still sufficient to provide experimentally observed backward-forward asymmetry, see Sec. VII.1.
V.3 The state
Quite reliable evidence for this state may be found in the data as a deep in the 7He spectrum at MeV and corresponding kink in the spectrum at MeV. However, the peak is sitting on a strong “background” of the and contributions, which are in general case comparable and not completely fixed by theoretical considerations. One of the fits to the data (attractive -wave potential) is shown in Fig. 10. The spectroscopic factors for the state are deduced by requesting a good fit quality in the energy range MeV are listed in the Table 1. The is expected to be a single-particle state with spectroscopic factor close to unity. Requesting the spectroscopic factor to be in the range we get limits on possible spectrum properties. The resonant energy positions outside the MeV interval are not acceptable. Considering the results of Sec. VII.1.2 as preferable explanation for the observed correlations, MeV seem to be a preferable value.
Attr. | Rep. | Attr. +FC | Rep. +FC | |
2.2 | 0.4 | 0.6 | 0.9 | 1.1 |
2.6 | 0.6 | 0.8 | 1.2 | 1.4 |
3.0 | 0.9 | 1.2 | 1.4 | 1.8 |
3.7 | 1.8 | 2.0 | 2.8 | 3.2 |
V.4 The coupled-channel state
Important aspects of the continuum dynamics beyond the single-channel potential approximation can be treated within the coupled-channel formalism. For our purposes we construct the couple-channel model is quite a schematic way. The coupled-channel Shrödinger equations look like
[…]
The channel potentials are obtained from - and - potentials in the folding model using the realistic 6He WFs, e.g. from [51]. The “SBB” - interaction [52] is used and simple effective - interaction [53]. The diagonal potentials for the 6He(g.s.)- channel and for the 6He()- channel are used practically “as is” (some fine-tuning can be applied). In contrast, the off-diagonal potential typically requires strong modification to provide sufficient coupling to get realistic mixing values for the 7He g.s.; it seem that this aspect of the 7He dynamics is beyond the folding approximation.
An example of the coupled-channel studies is shown in Fig. 8. The obtained in this model population probabilities have the following important features:
1.
The experimental 7He g.s. width is naturally well reproduced in this model.
2.
The second state is found at relatively low energy ( MeV).
3.
There is a strong deep in the population probability between the and states, which could be important for explanation of correlations in 7He decay, see Sec. VII.1.2.
4.
It can be found in Fig. 8 (b) that the off-shell extrapolation, which was quite safe for the single-channel potential model, is not working in the coupled-channel case.
VI Center-of-mass angular distributions
The center-of-mass angular distributions provide the standard basis for spin-parity identification in the direct reactions. For the reaction in inverse kinematics the recoil protons escaping in the relatively broad ( degrees) angular range in the backward direction in the lab system correspond to small c.m.s. angular range ( degrees). Therefore, the use of the reaction in our studies has both advantages and disadvantages.
The advantage for interpretation of our result is that population of states with high should be suppressed in this experiment. For example the only state, which is unambiguously identified in the excitation spectrum of 7He is the state at MeV, which decays with a large probability via the channel [1, 54]. However, population of this state requires compared to populating the and states. So, it is quite natural to expect that only the and resonant states of 7He are strongly populated in this experiment.
The disadvantage of the small angular range is that the cross section angular distribution is available in the narrow angular range and do not allow detailed comparison with reaction theory calculations to substantiate specific spin-parity prescription. Nevertheless, we are going to demonstrate that important conclusions are possible on the basis of the c.m.s. angular distributions derived in this experiment.
The experimental c.m.s. angular distributions for different excitation energy ranges of 7He are shown in Fig. 11. The data are corrected for registration efficiency of the setup via the MC procedure. The 7He ground state population cross section is found to be somewhat smaller than in Ref. [4]. The obtained angular distributions can be considered as independent of energy within the uncertainty of data.
The comparison of experimental data with the PWBA calculations is given in Fig. 12 (a). In its basic form, see Section IV, the PWBA model is not suitable for calculations of absolute cross sections. However, there is a standard phenomenological modification of the model: the peripheral Fermi-type cut-off for the “projectile” (deuteron) WF,
It can be seen in Fig. 12 (a) that both the 7He g.s. angular distributions of this work and of Ref. [3] are well reproduced with one parameter set fm and fm. It should be noted that spectroscopic factor for the 7He g.s. was taken the same in both calculations, while in the DWBA studies of [4] somewhat different values and were extracted for the 29 MeV data and for the 11.5 MeV data of Ref. [3].
The stability of the angular distributions, demonstrated in Fig. 11, is quite natural if the obtained data is strongly dominated by the same angular momentum transfer ( g.s., and excited states), in the whole excitation energy range MeV available in the experiment. The observation of the low-lying resonance in 7He was declared in [5], which should also exhibit itself in the angular distributions. Unfortunately, it is known, that specifically for the reactions the cross sections have highly analogous profiles. It can be seen in Fig. 12 (b) that to reliably distinguish the and contributions we should go to and need a data with a high statistical confidence. So, one of the ideas, how to prove or disprove the statement about the resonant state in 7He is to perform a dedicated experiment, in which a broad c.m.s. angular range is accessible.
VII Correlation studies for the 7He spectrum
[…]
In contrast, the longitudinal velocity of the heavy decay fragment (and, correspondingly, the projection of on the beam axis ) was measured for all events. For the events accepted by the setup of the experiment the equality
is precise within several percent. For this type of the data a broad excitation-energy range MeV is covered, but also the energy resolution is much lower ( keV FWHM). Due to low resolution of and for this data the correlations are available only in the form of backward-forward asymmetry in the 7He frame.
Let us first discuss the latter type of the correlations.
VII.1 Backward-forward asymmetry in the distribution of core longitudinal momentum
In our discussion of correlations the “forward” direction is the direction of and, correspondingly, or . The experimental backward-forward asymmetry function is defined as
see Fig. 13. Important feature of this distribution is rise from zero at MeV, then quite abrupt change of sign at about MeV, and then one more change of sign at MeV.
From Eq. (13) the backward-forward asymmetry is obtained as
Let us first demonstrate, how the observed backward-forward asymmetry can be qualitatively explained by a simple example of correlation pattern for narrow -wave states interference with flat -wave background, see Fig. 14. The population probabilities (12) and phase shifts for this example are provided by the standard R-matrix parameterization
We can find that the value is changing sign for the first time close to the resonance position (but not exactly at resonance), at the point where
since the contribution can be neglected here. The value is changing sign for the second time around the point
where the relative phase shift . The third time the sign is changed in analogy with Eq. (19) around the point
where the contribution can be neglected.
We should also point that even a very small, visually negligible, contribution of the -wave configuration could be sufficient to produce the typical observed asymmetry rate. The variation of the s-wave phase shift from attraction to repulsion leads only to insignificant variation of the whole correlation pattern, see and cases in Fig. 14.
The use of realistic profiles provided by potential model for the 7He -matrix gives qualitatively similar behavior, but do not provide quantitative agreement, see Fig. 15. There is a kink in the asymmetry function in between the and resonances, but the value is not falling towards zero: the effect of the state seem to be too small and it should be enlarged to achieve agreement with experiment. However, it can be found that contributions of the resonance are already on the upper limit admissible by the experimental MM spectrum, see Fig. 10. Alternatively we may inquire, whether the contribution of the state may be reduced at MeV. It can be seen in Fig. 15 (green double-dotted curve) that within the potential approach even the unrealistically strong variation of the g.s. properties does not lead to sufficiently strong variation of the high-energy tail of this state.
We have found three possible explanations for the data, as listed in the following Sections.
VII.1.1 Phase shift variation
The phase behavior of the -matrix elements is provided in PWBA as elastic phase shift in the 7He channel, see Eq. (11). However, as we have already mentioned in the Section IV, that this is not necessarily true in situation with some general reaction mechanism. It can be found that modifications of the T-matrix phase behavior
lead to results reasonably consistent with experimental observations, see Fig. 16. The first variant of modification in Eq. (22) looks more realistic, since the relative phase shift of -wave and -wave -matrices looks quite reasonable, say within DWBA, while the sizable relative phase shift of the and components is more complicated to justify.
The considered phase shift modification is beyond the simplistic PWBA discussed in this work. The question can be asked how realistic is such a modification, which we leave here as an opened question to reaction theory.
VII.1.2 Suppression of above MeV
Another possible explanation of the experimental asymmetry function, which naturally stems from Eq. (18) is connected with modification of the high-energy tail of the resonance. To check such possibility we have multiplied the population probabilities by Fermi function with energy parameter 1.3 MeV and width parameter 0.15 MeV. The results of calculations are shown in Fig. 17. There is good qualitative agreement at once and also a quantitative agreement with the data may be achieved by small variations of the phase shifts
Such a suppression of the high-energy tail of the resonance may find a natural explanation as a threshold effect in the 7He continuum. Indeed, the asymmetry function is changing sign at around MeV, which is suspiciously close to the 6He()+ channel threshold at MeV. To take this effect into account we need a coupled-channel calculations. Example of such calculations is shown in Fig. 8, where a strong suppression of the 6He()+ channel population is found around the 6He()+ threshold energy. We find this possibility the best candidate for reasonable explanation of experimental observations.
VII.1.3 The resonant state in 7He
Another possible explanation of the experimental asymmetry function, which naturally stems from Eq. (18) is modification of the continuum component, which can be connected with resonance somewhere in 7He continuum. The data interpretation in these terms was proposed in paper [5]. For this interpretation the R-matrix parameterization was used providing the phase shift
The first term here is R-matrix resonance phase shift parameterization for the -wave and the second term is just a hard-sphere-scattering phase shift. The calculation results with MeV, fm, and with -wave T-matrix calculated on-shell are shown in Fig. 18. The resonance width which can be formally associated with such parameterization is MeV. Qualitatively reasonable description of the observed asymmetry can be achieved.
This interpretation is based on two assumptions: (i) The off-shell interpolation of the elastic cross section induced by Eq. (24) was assumed to be trivial: the Authors of [5] followed Ref. [50] where the -wave off-shell interpolation is just constant; (ii) the parameters in the R-matrix expression Eq. (24) can be interpreted as actual resonance properties of the 7He continuum.
Concerning these assumptions we can comment the follows:
1.
The off-shell interpolation of the -wave T-matrix is not constant even in the simple potential model. It was shown by direct calculation in Fig. 9 that the off-shell modification of the T-matrix amplitudes can be as large as factor both up and down depending on particular properties of the 7He -wave interaction.
2.
A resonance in -wave (in contrast with virtual state) can be obtained only in the coupled-channel formulation of the scattering problem. For such formulation the -wave pole can arise due to coupling with resonant state in some, otherwise closed, channel. We performed the phenomenological studies within coupled-channel model trying to find the situation which can lead to phase shift behavior analogous to the one shown in Fig. 18. It was possible, but the deduced resonance properties of the continuum were found to be very different: MeV and MeV. This is the issue which is very familiar to practitioners of the R-matrix studies: for broad states the R-matrix parameters can be not easy to relate to physical observables. The -wave continuum just appeared especially weird in this sense.
3.
In the coupled-channel model the off-shell behavior of the T-matrix can be straightforwardly obtained. The results of such calculations are shown in Fig. 19. One may see that the off-shell behavior in such a model is complicated and qualitatively different from the on-shell behavior. So, if we really deal with the resonant state in -wave, the T-matrix should be only computed: there is no way to infer it from the on-shell (elastic) scattering T-matrix.
Thus the interpretation proposed in [5] is “phenomenologically possible” (as off-shell T-matrix parameterization describing the data) but can not be unambiguously related to “real physics” of the 7He continuum (described by the on-shell T-matrix).
VII.2 Correlations around the 7He g.s.
For this energy range the angular distribution for the 6He- relative motion can be reconstructed. We make it in the 7He c.m. frame with , where PWBA promise the most expressed correlation pattern. The data integrated over the MeV range is shown in Fig. 20 (c). For the 7He g.s. the description of the angular distribution can be confidently attributed to the interference. Then the whole distribution should be general parabolic shape
The distortions of individual terms from Eq. (25) by the experimental setup were studied by using the MC procedure and we may see in Fig. 20 (c) that these distortions are modest.
For the pure the PWBA model Eq. (13) predicts the expressed parabolic profile with “hill-to-valley” ratio equal 4, see Fig. 20 (a,b,d) black solid curve. Let’s discuss meaning of this fact and whether some additional information could be extracted from this distribution.
Some asymmetry can be found in the distribution Fig. 20 (c). However, this asymmetry is actually difficult to relate to physics of the experiment. The integrated asymmetry effect connected with - interference at the 7He g.s. is quite small because asymmetry is changing sign around the resonant energy and the observed small value is result of fine compensation of large values from below and from above of the resonance. Fig. 21 shows how this effect depends on the integration range. It is clear that in any experiment, where the energy resolution is lower than the width of the 7He g.s. such an asymmetry actually shows the integration energy range which we can afford because of resolution and statistics. It can be also found from Fig. 21 that there is large sensitivity of the integrated asymmetry value on the phase convention. Potentially, this property can be used to determine the phase convention experimentally. At the moment the quality of the data is not sufficient (the error bars shown in this plot are pure statistical, the actual uncertainty is larger). The dependence of the measured asymmetry on the experimental resolution is further elaborated in Fig. 22. Again we can find that potentially, the asymmetry function can be used to determine the phase convention experimentally. However, at the moment the comparison of Fig. 22 can lead only to qualitative conclusions as the c.m. spectrum has good enough resolution, but low-statistics data with large error bars, while the MM spectrum reliable statistically suffer from large effect of energy resolution.
Another aspect of the angular distributions in Fig. 20 is connected with - interference and alignment properties.
The - interference averaged over the energy integration range is illustrated in Fig. 20 (a). The effect is small, whatever relative - phase is assumed. Actually we expect this phase close to at the resonance position, and therefore small interference effects anyhow. So, here no sensitivity is found and no information can be extracted.
Situation with alignment is more interesting. Among configurations , , and only the one may be spin-aligned. It is convenient to discuss the alignment in terms of the alignment coefficient
It is clear that for pure configuration, the alignment lead to the isotropic angular distribution (this is case of “statistical” population of magnetic substates), see Fig. 20 (b), green dotted curve. The complete “equatorial” alignment with leads to an expressed parabolic concave profile with “baseline” equal to 0.25, see Fig. 20 (b), solid black curve. The “polar” alignment with leads to parabolic convex profiles. The latter profile is clearly in a strong disagreement with the data. The experimental data Fig. 20 (d) clearly requires pure or practically pure equatorial alignment with . Pure equatorial alignment is trivially predicted in PWBA or similar models, which can be qualitatively attributed to the single-step (or “single-pole”) reaction mechanism. So, the observed angular distribution confirming the near-perfect equatorial alignment can be seen as strong support of the applicability of the used theoretical approach.
So, to finalize this point, the observed angular correlations for the 7He g.s. decay provide a strong confirmation for a “single-pole” single-step reaction mechanism leading to population of 7He ground state. We have shown that the data of this kind, but higher quality can be used to establish solidly the phase convention for - interference in proximity of 7He ground state.
VIII Outlook
The results obtained in this work show nice prospects of the discussed correlation method for detailed and precise studies of the 7He continuum. The following issues may be addressed in the forthcoming studies of the system.
1.
The use of the “combined mass” approach allows to drastically improve the energy resolution of the 7He spectrum. Aiming experiment with high statistics we may pretend to derive the 7He g.s. width with precision keV. The 7He g.s. is well known to be not a single particle state with the dominant + configuration with spectroscopic factors varying in different studies in the range (e.g. [6, 7]). The improved width data would allow to elaborate this question.
2.
The variants of explanation for the asymmetry function formulated in Sections VII.1.1, VII.1.2, and VII.1.3 are to certain extent connected with the limited character of information provided by asymmetry. This drawback could be overcome in a more sophisticated experiment, where high-quality complete correlation information will be available in a broad energy range. The illustration of the predicted complete correlation pattern is given in Fig. 23. The important aspect of information which is missing in the current data is transition convex-concave for the parabolic component of the 6He- angular distribution. The corresponding part of the cross section is provided in PWBA as
(26) and it is clear that the coefficient at the term is changing sign in the points providing different information compared to the asymmetry coefficient Eq. (18).
3.
The limitations on the properties of the resonance are obtained in this work are not very restrictive. There are two main reasons. (i) Unresolved contribution of the - inelastic channel in the experiment. This issue can be easily resolved in a more “accurate” design of the experimental setup providing the clear identification of 4He and 6He. (ii) Uncertainty in the population of the -wave continuum. This issue is more complicated as main qualitative properties of the correlations retain, for example, for 10-fold (even 50-fold) reduction of the -wave continuum population compared to the practically maximal possible contribution used in the current analysis. However, potentially it can also be resolved if the precise data of the type shown in Fig. 23 is available.
IX Conclusions
The 7He continuum states were studied in the 2H(6He,1H)7He reaction at 29 MeV. As compared to previous studies in the same reaction [2, 3] this experiment provides higher statistics, higher resolution, and larger excitation energy coverage. The details of the analysis procedure are somewhat different from those published in [4, 5] and there are important differences in interpretation of the data. The interpretation of the data in this work is based on extensive PWBA plus coupled-channel studies. They are indicating extreme importance of the careful treatment of the off-shell effects for understanding of the correlation patterns induced in 7He by the reaction.
The population cross section for the 7He g.s. is found to be consistent with the data [3] obtained at 11.5 MeV. The c.m. angular distributions for the reaction are practically independent on the 7He excitation energy, supporting the same reaction mechanism and the same dominant up to MeV.
The 7He g.s. properties are established as MeV and keV. Angular distribution for the 6He- decay of this state can be explained by a strong spin alignment induced by a reaction mechanism. The equatorial spin alignment is a natural feature of the single-step single-pole direct reaction models including PWBA. Thus this fact is a solid confirmation of the robustness of our theoretical considerations.
The correlation information for the higher-lying 7He excitations is available as backward-forward asymmetry for the 6He- decay in the 7He frame aligned with transferred momentum . The asymmetry function has an expressed profile with three sign-changing points (the first of them is very likely from the data and two others are reliably observed). Such a behavior of this function may be explained by using quite restrictive assumptions and only three such explanations were found:
1.
Strong phase variation (compared to PWBA) due to more complicated reaction mechanism. Possibility of such an interpretation could be confirmed or dismissed by the further advanced reaction theory studies.
2.
Suppression of continuum above MeV. This interpretation gets support from the coupled channel calculations as a threshold effect in proximity to the 6He()+ channel threshold.
3.
The existence of the resonant state in 7He was declared in Ref. [5] with , MeV. This interpretation is in contradiction with the coupled-channel calculations, which predict extremely strong off-shell effects in this case, providing qualitatively different picture after off-shell correction. No special need for the resonant state is found in the data analysis of this work.
In all the above scenarios the first excited state of 7He is (quite naturally) state with peak positions confined to a range MeV with “preferred” value around 2.6 MeV. There is indication on the second state in the data with MeV and the lower resonance energy limit MeV.
In this work it was demonstrated that the discussed correlation method has the potential to solve the listed burning questions of the 7He continuum, however it is clear that prospective experiments should have higher statistics, higher energy resolution and unambiguous identification of decay channels.
Acknowledgements.
We are grateful to Prof. A.S. Fomichev for important discussions. The research was supported in part in the framework of scientific program of the Russian National Center for Physics and Mathematics, topic number 6 “Nuclear and radiation physics” (2023–2025 stage). We acknowledge the interest and support of this activity from Profs. Yu.Ts. Oganessian, S.N. Dmitriev and S.I. Sidorchuk.References
- Korsheninnikov et al. [1999] A. A. Korsheninnikov, M. S. Golovkov, A. Ozawa, E. A. Kuzmin, E. Y. Nikolskii, K. Yoshida, B. G. Novatskii, A. A. Ogloblin, I. Tanihata, Z. Fulop, K. Kusaka, K. Morimoto, H. Otsu, H. Petrascu, and F. Tokanai, Phys. Rev. Lett. 82, 3581 (1999).
- Golovkov et al. [2001] M. S. Golovkov, A. A. Korsheninnikov, I. Tanihata, D. D. Bogdanov, M. L. Chelnokov, A. S. Fomichev, V. A. Gorshkov, Y. T. Oganessian, A. M. Rodin, S. I. Sidorchuk, S. V. Stepantsov, G. M. Ter-Akopian, R. Wolski, W. Mittig, P. Roussel-Chomaz, H. Savajols, E. A. Kuzmin, E. Y. Nikolskii, B. G. Novatskii, and A. A. Ogloblin, Physics of Atomic Nuclei 64, 1244 (2001).
- Wuosmaa et al. [2005] A. H. Wuosmaa, K. E. Rehm, J. P. Greene, D. J. Henderson, R. V. F. Janssens, C. L. Jiang, L. Jisonna, E. F. Moore, R. C. Pardo, M. Paul, D. Peterson, S. C. Pieper, G. Savard, J. P. Schiffer, R. E. Segel, S. Sinha, X. Tang, and R. B. Wiringa, Phys. Rev. C 72, 061301 (2005).
- Bezbakh et al. [2024] A. A. Bezbakh, M. S. Golovkov, A. S. Denikin, R. Wolski, S. G. Belogurov, D. Biare, V. Chudoba, A. S. Fomichev, E. M. Gazeeva, A. V. Gorshkov, G. Kaminski, B. R. Khamidullin, M. Khirk, S. A. Krupko, B. Mauyey, I. A. Muzalevskii, W. Piatek, A. M. Quynh, S. I. Sidorchuk, R. S. Slepnev, A. Swiercz, G. M. Ter-Akopian, and B. Zalewski, International Journal of Modern Physics E 33, 2450002 (2024).
- Golovkov et al. [2024] M. S. Golovkov, A. A. Bezbakh, A. S. Denikin, R. Wolski, S. G. Belogurov, D. Biare, V. Chudoba, E. M. Gazeeva, A. V. Gorshkov, G. Kaminski, B. R. Khamidullin, S. A. Krupko, B. Mauyey, I. A. Muzalevskii, W. Piatek, A. M. Quynh, R. S. Slepnev, A. Swiercz, P. Szymkiewicz, and B. Zalewski, Phys. Rev. C 109, L061602 (2024).
- Renzi et al. [2016] F. Renzi, R. Raabe, G. Randisi, D. Smirnov, C. Angulo, J. Cabrera, E. Casarejos, T. Keutgen, A. Ninane, J. L. Charvet, A. Gillibert, V. Lapoux, L. Nalpas, A. Obertelli, F. Skaza, J. L. Sida, N. A. Orr, S. I. Sidorchuk, R. Wolski, M. J. G. Borge, and D. Escrig, Phys. Rev. C 94, 024619 (2016).
- Fortune [2018] H. T. Fortune, Eur. Phys. J. A 54, 51 (2018).
- Markenroth et al. [2001] K. Markenroth, M. Meister, B. Eberlein, D. Aleksandrov, T. Aumann, L. Axelsson, T. Baumann, M. Borge, L. Chulkov, W. Dostal, T. Elze, H. Emling, H. Geissel, A. Grunschloss, M. Hellstrom, J. Holeczek, B. Jonson, J. Kratz, R. Kulessa, A. Leistenschneider, I. Mukha, G. Munzenberg, F. Nickel, T. Nilsson, G. Nyman, M. Pfutzner, V. Pribora, A. Richter, K. Riisager, C. Scheidenberger, G. Schrieder, H. Simon, J. Stroth, O. Tengblad, and M. Zhukov, Nuclear Physics A 679, 462 (2001).
- Meister et al. [2002] M. Meister, K. Markenroth, D. Aleksandrov, T. Aumann, T. Baumann, M. Borge, L. Chulkov, D. Cortina-Gil, B. Eberlein, T. Elze, H. Emling, H. Geissel, M. Hellstrom, B. Jonson, J. Kratz, R. Kulessa, A. Leistenschneider, I. Mukha, G. Munzenberg, F. Nickel, T. Nilsson, G. Nyman, M. Pfutzner, V. Pribora, A. Richter, K. Riisager, C. Scheidenberger, G. Schrieder, H. Simon, O. Tengblad, and M. Zhukov, Nuclear Physics A 700, 3 (2002).
- Halderson [2004] D. Halderson, Phys. Rev. C 70, 041603 (2004).
- Skaza et al. [2006] F. Skaza, V. Lapoux, N. Keeley, N. Alamanos, E. C. Pollacco, F. Auger, A. Drouart, A. Gillibert, D. Beaumel, E. Becheva, Y. Blumenfeld, F. Delaunay, L. Giot, K. W. Kemper, L. Nalpas, A. Obertelli, A. Pakou, R. Raabe, P. Roussel-Chomaz, J.-L. Sida, J.-A. Scarpaci, S. Stepantsov, and R. Wolski, Phys. Rev. C 73, 044301 (2006).
- Canton et al. [2006] L. Canton, G. Pisent, K. Amos, S. Karataglidis, J. P. Svenne, and D. v. d. Knijff, Phys. Rev. C 74, 064605 (2006).
- Ryezayeva et al. [2006] N. Ryezayeva, C. Bäumer, A. van den Berg, L. Chulkov, D. Frekers, D. De Frenne, E.-W. Grewe, P. Haefner, E. Jacobs, H. Johanson, Y. Kalmykov, A. Negret, P. von Neumann-Cosel, L. Popescu, S. Rakers, A. Richter, G. Schrieder, A. Shevchenko, H. Simon, and H. Wörtche, Physics Letters B 639, 623 (2006).
- Myo et al. [2007] T. Myo, K. Kato, and K. Ikeda, Phys. Rev. C 76, 054309 (2007).
- Arai and Aoyama [2009] K. Arai and S. Aoyama, Phys. Rev. C 80, 027301 (2009).
- Poppelier et al. [1985] N. Poppelier, L. Wood, and P. Glaudemans, Physics Letters B 157, 120 (1985).
- Wolters et al. [1990] A. A. Wolters, A. G. M. van Hees, and P. W. M. Glaudemans, Phys. Rev. C 42, 2062 (1990).
- Wurzer and Hofman [1996] J. Wurzer and H. M. Hofman, Phys. Rev. C 55, 688 (1996), rGM calculations.
- Navrátil and Barrett [1998] P. Navrátil and B. R. Barrett, Phys. Rev. C 57, 3119 (1998).
- Bohlen et al. [2001] H. G. Bohlen, R. Kalpakchieva, A. Blažević, B. Gebauer, T. N. Massey, W. von Oertzen, and S. Thummerer, Phys. Rev. C 64, 024312 (2001).
- Mittig et al. [2005] W. Mittig, C. E. Demonchy, H. Wang, P. Roussel-Chomaz, B. Jurado, M. Gelin, H. Savajols, A. Fomichev, A. Rodin, A. Gillibert, A. Obertelli, M. D. Cortina-Gil, M. Caama no, M. Chartier, and R. Wolski, The European Physical Journal A - Hadrons and Nuclei 25, 263 (2005).
- Aksyutina et al. [2009] Y. Aksyutina, H. Johansson, T. Aumann, K. Boretzky, M. Borge, A. Chatillon, L. Chulkov, D. Cortina-Gil, U. D. Pramanik, H. Emling, C. Forssén, H. Fynbo, H. Geissel, G. Ickert, B. Jonson, R. Kulessa, C. Langer, M. Lantz, T. LeBleis, A. Lindahl, K. Mahata, M. Meister, G. Münzenberg, T. Nilsson, G. Nyman, R. Palit, S. Paschalis, W. Prokopowicz, R. Reifarth, A. Richter, K. Riisager, G. Schrieder, H. Simon, K. Sümmerer, O. Tengblad, H. Weick, and M. Zhukov, Physics Letters B 679, 191 (2009).
- Pieper et al. [2002] S. C. Pieper, K. Varga, and R. B. Wiringa, Phys. Rev. C 66, 044310 (2002).
- Baroni et al. [2013] S. Baroni, P. Navrátil, and S. Quaglioni, Phys. Rev. C 87, 034326 (2013).
- Myo et al. [2009] T. Myo, R. Ando, and K. Katō, Phys. Rev. C 80, 014315 (2009).
- Rodkin and Tchuvil’sky [2021] D. M. Rodkin and Y. M. Tchuvil’sky, Phys. Rev. C 104, 044323 (2021).
- Chen et al. [2001] L. Chen, B. Blank, B. Brown, M. Chartier, A. Galonsky, P. Hansen, and M. Thoennessen, Physics Letters B 505, 21 (2001).
- Al Kalanee et al. [2013] T. Al Kalanee, J. Gibelin, P. Roussel-Chomaz, N. Keeley, D. Beaumel, Y. Blumenfeld, B. Fernández-Domínguez, C. Force, L. Gaudefroy, A. Gillibert, J. Guillot, H. Iwasaki, S. Krupko, V. Lapoux, W. Mittig, X. Mougeot, L. Nalpas, E. Pollacco, K. Rusek, T. Roger, H. Savajols, N. de Séréville, S. Sidorchuk, D. Suzuki, I. Strojek, and N. A. Orr, Phys. Rev. C 88, 034301 (2013).
- Golovkov et al. [2007] M. S. Golovkov, L. V. Grigorenko, A. S. Fomichev, A. V. Gorshkov, V. A. Gorshkov, S. A. Krupko, Y. T. Oganessian, A. M. Rodin, S. I. Sidorchuk, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, R. Wolski, A. A. Korsheninnikov, E. A. Kuzmin, E. Y. Nikolskii, B. G. Novatskii, D. N. Stepanov, P. Roussel-Chomaz, and W. Mittig, Phys. Rev. C 76, 021605(R) (2007).
- Fortier et al. [2007] S. Fortier, E. Tryggestad, E. Rich, D. Beaumel, E. Becheva, Y. Blumenfeld, F. Delaunay, A. Drouart, A. Fomichev, N. Frascaria, S. Gales, L. Gaudefroy, A. Gillibert, J. Guillot, F. Hammache, K. W. Kemper, E. Khan, V. Lapoux, V. Lima, L. Nalpas, A. Obertelli, E. C. Pollacco, F. Skaza, U. D. Pramanik, P. Roussel-Chomaz, D. Santonocito, J. A. Scarpaci, O. Sorlin, S. V. Stepantsov, G. M. Ter Akopian, and R. Wolski, AIP Conference Proceedings 912, 3 (2007), https://aip.scitation.org/doi/pdf/10.1063/1.2746575 .
- Jaganathen et al. [2017] Y. Jaganathen, R. M. I. Betan, N. Michel, W. Nazarewicz, and M. Płoszajczak, Phys. Rev. C 96, 054316 (2017).
- Myo and Kato [2020] T. Myo and K. Kato, Progress of Theoretical and Experimental Physics 2020, 12A101 (2020), https://academic.oup.com/ptep/article-pdf/2020/12/12A101/35612142/ptaa101.pdf .
- Fossez et al. [2018] K. Fossez, J. Rotureau, and W. Nazarewicz, Phys. Rev. C 98, 061302 (2018).
- Mazur et al. [2022] I. A. Mazur, I. J. Shin, Y. Kim, A. I. Mazur, A. M. Shirokov, P. Maris, and J. P. Vary, Phys. Rev. C 106, 064320 (2022).
- Fomichev et al. [2018] A. S. Fomichev, L. V. Grigorenko, S. A. Krupko, S. V. Stepantsov, and G. M. Ter-Akopian, The European Physical Journal A 54, 97 (2018).
- Bezbakh et al. [2020] A. A. Bezbakh, S. G. Belogurov, D. Biare, V. Chudoba, A. S. Fomichev, E. M. Gazeeva, M. S. Golovkov, A. V. Gorshkov, G. Kaminski, S. A. Krupko, B. Mauyey, I. A. Muzalevskii, E. Y. Nikolskii, Y. L. Parfenova, W. Piatek, A. M. Quynh, A. Serikov, S. I. Sidorchuk, P. G. Sharov, R. S. Slepnev, S. V. Stepantsov, A. Swiercz, P. Szymkiewicz, G. M. Ter-Akopian, R. Wolski, and B. Zalewski, Bulletin of the Russian Academy of Sciences: Physics 84, 491 (2020).
- Bezbakh et al. [2018] A. A. Bezbakh, S. G. Belogurov, R. Wolski, E. M. Gazeeva, M. S. Golovkov, A. V. Gorshkov, G. Kaminski, M. Y. Kozlov, S. A. Krupko, I. A. Muzalevky, E. Y. Nikolskii, E. V. Ovcharenko, R. S. Slepnev, G. M. Ter-Akopian, A. S. Fomichev, P. G. S. V. Chudoba, and V. N. Schetinin, Instruments and Experimental Techniques 61, 631 (2018).
- Sharov et al. [2017] P. G. Sharov, A. S. Fomichev, A. A. Bezbakh, V. Chudoba, I. A. Egorova, M. S. Golovkov, T. A. Golubkova, A. V. Gorshkov, L. V. Grigorenko, G. Kaminski, A. G. Knyazev, S. A. Krupko, M. Mentel, E. Y. Nikolskii, Y. L. Parfenova, P. Pluchinski, S. A. Rymzhanova, S. I. Sidorchuk, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, and R. Wolski, Phys. Rev. C 96, 025807 (2017).
- Golovkov et al. [2004] M. S. Golovkov, L. V. Grigorenko, A. S. Fomichev, Y. T. Oganessian, Y. I. Orlov, A. M. Rodin, S. I. Sidorchuk, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, and R. Wolski, Phys. Lett. B 588, 163 (2004).
- Golovkov et al. [2005] M. S. Golovkov, L. V. Grigorenko, A. S. Fomichev, S. A. Krupko, Y. T. Oganessian, A. M. Rodin, S. I. Sidorchuk, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, R. Wolski, M. G. Itkis, A. S. Denikin, A. A. Bogatchev, N. A. Kondratiev, E. M. Kozulin, A. A. Korsheninnikov, E. Y. Nikolskii, P. Roussel-Chomaz, W. Mittig, R. Palit, V. Bouchat, V. Kinnard, T. Materna, F. Hanappe, O. Dorvaux, L. Stuttgé, C. Angulo, V. Lapoux, R. Raabe, L. Nalpas, A. A. Yukhimchuk, V. V. Perevozchikov, Y. I. Vinogradov, S. K. Grishechkin, and S. V. Zlatoustovskiy, Phys. Rev. C 72, 064612 (2005).
- Sidorchuk et al. [2010] S. I. Sidorchuk, A. S. Fomichev, M. S. Golovkov, L. V. Grigorenko, V. A. Gorshkov, A. V. Gorshkov, S. A. Krupko, Y. T. Oganessian, A. M. Rodin, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, and R. Wolski, Nucl. Phys. A 840, 1 (2010).
- Sidorchuk et al. [2012] S. I. Sidorchuk, A. A. Bezbakh, V. Chudoba, I. A. Egorova, A. S. Fomichev, M. S. Golovkov, A. V. Gorshkov, V. A. Gorshkov, L. V. Grigorenko, P. Jalůvková, G. Kaminski, S. A. Krupko, E. A. Kuzmin, E. Y. Nikolskii, Y. T. Oganessian, Y. L. Parfenova, P. G. Sharov, R. S. Slepnev, S. V. Stepantsov, G. M. Ter-Akopian, R. Wolski, A. A. Yukhimchuk, S. V. Filchagin, A. A. Kirdyashkin, I. P. Maksimkin, and O. P. Vikhlyantsev, Phys. Rev. Lett. 108, 202502 (2012).
- Tilley et al. [2002] D. Tilley, C. Cheves, J. Godwin, G. Hale, H. Hofmann, J. Kelley, C. Sheu, and H. Weller, Nucl. Phys. A708, 3 (2002).
- Beck et al. [2007] F. Beck, D. Frekers, P. von Neumann-Cosel, A. Richter, N. Ryezayeva, and I. Thompson, Physics Letters B 645, 128 (2007).
- Denby et al. [2008] D. H. Denby, P. A. DeYoung, T. Baumann, D. Bazin, E. Breitbach, J. Brown, N. Frank, A. Gade, C. C. Hall, J. Hinnefeld, C. R. Hoffman, R. Howes, R. A. Jenson, B. Luther, S. M. Mosby, C. W. Olson, W. A. Peters, A. Schiller, A. Spyrou, and M. Thoennessen, Phys. Rev. C 78, 044303 (2008).
- Cao et al. [2012] Z. Cao, Y. Ye, J. Xiao, L. Lv, D. Jiang, T. Zheng, H. Hua, Z. Li, X. Li, Y. Ge, J. Lou, R. Qiao, Q. Li, H. You, R. Chen, D. Pang, H. Sakurai, H. Otsu, M. Nishimura, S. Sakaguchi, H. Baba, Y. Togano, K. Yoneda, C. Li, S. Wang, H. Wang, K. Li, T. Nakamura, Y. Nakayama, Y. Kondo, S. Deguchi, Y. Satou, and K. Tshoo, Physics Letters B 707, 46 (2012).
- Sitenko [1959] A. Sitenko, Sov. Phys. Usp. 2, 195 (1959).
- Shapiro [1961] I. Shapiro, Nucl. Phys. 28, 244 (1961).
- Fröbrich and Lipperheide [1996] P. Fröbrich and R. Lipperheide, Theory of Nuclear Reactions (Clarendon Press, Oxford Studies in Nuclear Physics, 1996).
- Fuchs et al. [1972] H. Fuchs, H. Homeyer, H. Oeschler, R. Lipperheide, and K. Möhring, Nuclear Physics A 196, 286 (1972).
- Grigorenko et al. [2020] L. V. Grigorenko, N. B. Shulgina, and M. V. Zhukov, Phys. Rev. C 102, 014611 (2020).
- Sack et al. [1954] S. Sack, L. C. Biedenharn, and G. Breit, Phys. Rev. 93, 321 (1954).
- Cooper and Eisenberg [1968] B. S. Cooper and J. Eisenberg, Nuclear Physics A 114, 184 (1968).
- Wuosmaa et al. [2008] A. H. Wuosmaa, J. P. Schiffer, K. E. Rehm, J. P. Greene, D. J. Henderson, R. V. F. Janssens, C. L. Jiang, L. Jisonna, J. C. Lighthall, S. T. Marley, E. F. Moore, R. C. Pardo, N. Patel, M. Paul, D. Peterson, S. C. Pieper, G. Savard, R. E. Segel, R. H. Siemssen, X. D. Tang, and R. B. Wiringa, Phys. Rev. C 78, 041302 (2008).